### 04 Jan 2024 v2

# Remembering Nino Boccara (1931–2018)

Remembering Nino Boccara (1931–2018)
Henryk Fukś
1 Brief biography
Nino (Samson) Boccara was born in Tunis on May 30, 1931, but when he was still a child, his family moved to France. He attended one of the most reputable schools in Paris, the École alsacienne, and in 1952 entered the highly competitive École supérieure de physique et de chimie industrielles de la ville de Paris (ESPCI Paris), one of the top engineering schools in France, today famous for having five Nobel prize winners among its former professors and alumni. He graduated from ESPCI in 1956 as a member of the 71st graduating class with the diploma of engineer and then continued as a graduate student at the University of Paris (Sorbonne) under the direction of Jean Laval (1900–1980), professor of theoretical physics at the Collège de France from 1950 to 1970.
In December 1961, he defended the doctoral thesis entitled Étude de l’effet Compton et des ondes élastiques dans les cristaux de sylvine (Study of the Compton effect and elastic waves in sylvite crystals) receiving the degree of “Docteur ès sciences physiques” (PhD). After the doctorate he worked for the next two years at the Laboratory of Theoretical Physics of Collège de France, and in 1964 secured a position at the Centre d’Etudes Nucléaires de Saclay, research centre belonging to the Atomic Energy Commision (CEA), today known as CEA Paris-Saclay. At first, he worked in the unit called Solid State Physics and Magnetic Resonace Service, and later, after reorganization which happened around 1989/1990, in DRECAM/SPEC (Research Division on Atoms, Molecules and Condensed Matter/Service of Condensed Matter Physics). He remained affiliated with Saclay until retirement.
In addition to his position at Saclay, he held appointments at several other institutions. In 1977 he became a professor of mathematics at ESPCI, where he taught various courses in mathematics for many years, continuing well into late 1990’s. He held the title of directeur de recherche (senior scientist) at the CNRS, Centre national de la recherche scientifique (National Centre for Scientific Research). Furthermore, for a number of years he served as a director of Centre de Physique at Les Houches, where throughout 1980’s and in early 90’s he organized a series of workshops and edited proceedings books [boccara1981symmetries, boccara85physics, leger87polycyclic, maret1986biophysical, mlb87, lelay89semiconductor, richard1988elementary, jullien1988universalities, manneville1989cellular, polian1989simple, cb90, bgmp93]. In 1989, he joined the Physics Department of the University of Illinois at Chicago as a full professor, a position which he also held until retirement. He had some kind of arrangement with both UIC and Saclay which allowed him to spend half of the year in Paris and the other half in Chicago.
He retired from UIC in 2002, and around that time also from other positions he held in France. He continued research and publishing activities well beyond this date, with the last research paper published in 2010. Nino Boccara died in Paris on December 16, 2018 at the age of 87.
Figure 1: Top: the 71st graduating class of ESPCI, 1957 (from https://www.espci.org/fr/trombi/medium/71.jpg). N.B. in the first row standing, third from the right. Bottom: Front page of the PhD thesis of N.B. from 1961. Reproduced from https://inis.iaea.org/collection/NCLCollectionStore/_Public/54/034/54034400.pdf.
2 Personal recollections
I met Nino Boccara for the first time relatively late in his career, in 1993, when I started my graduate studies at the University of Illinois at Chicago. After passing the preliminary exam, I was looking for an advisor, and then I read the departamental brochure describing research of all physics professors. In this brochure, his research was described under the heading ,,Discrete Dynamical Systems” and the text below the heading was as follows:
Most models in population dynamics are usually formulated in terms of differential equations or random processes. These models cannot take into account the local character of the interactions, which is, on the contrary, included in automata network models. In collaboration with graduate students at UIC and other collaborators in France, Professor Boccara has build up various models in epidemiology and ecology. More recently, he has submitted a report to the Illinois Criminal Justice Authority on street gang growth control.
This was so very different than the research one would expect to find at the physics department that I became very intrigued and decided to talk to him. Shortly afterwards, after two or three long meetings, he officially became my advisor. For the next few years I worked with him on various aspects of the theory and applications of cellular automata. We continued collaboration also after my PhD and published a number of joint papers, with the last one in 2006.
Nino Boccara taught at UIC various physics courses at both undergraduate and graduate level. In the Fall 1994, I took his Statistical Mechanics course, PHYS 561, which I enjoyed tremendously, although I remember it as being very hard, especially the final exam. He incorporated cellular automata (CA) into the course and showed us how some methods of statistical physics could be used to analyze dynamics of CA. This is when I learned about the generalization of the mean-field approximation known as the local structure theory, developed by Howard Gutowitz in late 80’s.
When I came to UIC, he had several graduate students. His first student, Kyeong Taik Cheong, already defended his PhD thesis in 1992, but Qi Zeng, Mark Oram and Servie Geurten either continued or arrived during my stay. Franco Bagnoli also started his PhD under Nino’s direction at approximately the same time, at University of Paris VI, where Nino had one of his multiple affiliations.
Nino was giving his students a lot of freedom to choose research topics and did not really press too hard for results. His students, therefore, had to be able to work independently and needed a measure of self-discipline in order to make progress in thesis research. Because he was traveling a lot and spending time at different institutions, this also meant that much of the interaction with him had to be done remotely. Here is how Franco Bagnoli remembers working with Nino:
I do not remember exactly when I met Nino for the first time, probably at the workshop “Complexity and Evolution” in Les Houches [^1] Workshop which took place at Les Houches in March 6-15, 1990. in 1990. We also met on other occasions, likely in Turin. Anyhow, he was already one of the “big” names in the field of cellular automata, the subject of my master thesis (1989), and I was familiar with many of his papers. In 1991, after having spent some time in Geneva (Switzerland) working with Michel Droz, I was hired by the University of Florence as a technician, the system manager of the Department of Applied Mathematics, which curiously at that time had no network, no servers and just a couple of PCs. I won the competition for that position because I used to work in a computer company during my master studies, and because the examination was organized by mathematicians, with questions about functional analysis and linear algebra mixed with computation.
After having built from scratch the network and the Internet services of the department, I decided that after all I preferred doing physics rather than informatics for a living, so I decided to get a PhD. I was hired full-time (although with a lot of freedom in my schedule), and therefore I looked for a PhD abroad, since at that time in Italy it was not possible to pursue a PhD program without a grant (and thus without abandoning my job). I contacted some researchers in France and Germany and finally I got a response from Nino, who agreed to be the director of my PhD.
At that time he was working partially in ESPCI in Paris and in the University of Illinois at Chicago, but he also had some affiliation or connection with the University Paris VI (Pierre et Marie Curie), and CEA Paris-Saclay. So he asked me which of these institutions I preferred for my PhD. Since ESPCI was mainly an engineering school, I chose Paris VI which was also the easiest to reach from Florence, being in the center of Paris. I did not want to leave my job, so I had to use my free time and days off. In the end we have never worked in Paris VI, we always met in Saclay when I traveled to Paris, and a couple of times in Chicago. He also owned a house in central Italy, in Magione, near Trasimeno lake, and we met there once. For most of the time, we collaborated by email (well before the existence of skype or other chat systems), and I discussed my thesis in 1997 in Saclay.
I have to say that relatively little of my thesis originated from the collaboration with Nino. We published only one paper together, while I collected about 10 articles for my PhD. Afterwards, we met from time to time on various occasions, and he participated as a teacher in a school that I co-organized in Turin, in 2000. So, although Nino was always a brilliant mind, I profited more from him before and after my PhD than during it, essentially because we had few occasions to interact, he was always travelling and my mind was split between computer science and theoretical physics. Anyhow, I finally managed to become a researcher (in 2001) and then a professor of physics (with computer science as a hobby), so I can say that Nino helped me in reaching my goals.
In the mid-90’s computers and internet technology were rapidly evolving and Nino was an enthusiastic adopter of all new technologies. He was a big fan of Macintosh computers and had in his office both desktop and laptop versions of Mac. Since I did not share his enthusiasm for Apple, he sometimes tried to convince me how much better Macs are compared to PCs, and usually I had to concede. This was also the time when many other operating systems competed: VM, Vax, OS2, and many flavours of Unix. Nino tried to stay on top of all these things and we experimented with running cellular automata simulations on various servers which Physics Department had. I should add here that Nino was also a serious Mathematica user, incorporated symbolic algebra into his courses and later wrote a book on Mathematica programming [boccara2007essentials].
He possessed tremendous and very broad knowledge of diverse areas of mathematics. For his students at ESPCI he wrote a series of textbooks ,,Mathematics For Engineers”, for courses in probability, integration, analytic functions and theory of distributions [boccara1995probabilites, boccara1995integration, boccara1996fonctions, boccara1997distributions]. His most popular mathematics textbook was Functional analysis: an introduction for physicists, first published in French [boccara1984analyse] and then in English by Elsevier [boccara1990functional]. Speaking of books, he authored 11 books, but has never became a skilled typist. A fellow graduate student at UIC once laughingly remarked that Nino wrote all these books using one finger only…
Although, as Franco Bagnoli remarked above, Nino was frequently traveling and his students sometimes had to be patient to meet him in person, when he was around he always had time for students and colleagues. He was a man of great culture, spoke fluently three languages, widely traveled, and had a great range of interests. I always admired his erudition and his French-Italian sense of humor. He is sorely missed by everyone fortunate enough to have known him.
Below, I will discuss the works of Nino Boccara in more or less chronological order. I will place somewhat more emphasis on the period after 1989, primarily because I am more familiar with his research during the later stage of his career.
3 Statistical physics, 1960–1989
It is very difficult to describe the vast research output of Nino Boccara in one short note. I will, therefore, only describe some highlights here, without any pretension of being exhaustive. For the rest, I refer the reader to the the bibliography section included at the end. I believe that the bibliography is complete for the period 1989–2010, but for earlier years I probably missed some publications in conference proceedings or in some less known journals.
As already mentioned, his doctoral thesis was on Compton scattering, and for a while he continued this line of work [nb1960c, nb1960b, nb1960a, WOS:A1962WS83000003, WOS:A1962WS83600009, Wos:A19639216a00006, Wos:A19639257a00007, 1963SSCom...1..168B]. In mid 1960’s his interests shifted toward the theory of phase transitions. At the beginning, he studied phase transitions induced by deformation of crystals. His most cited paper from that time, published in Annals of Physics in 1968 [1968AnPhy..47...40B], examines phase transitions accompanied by a change in crystal structure and demonstrates how to use symmetry arguments (Landau’s theory) to determine all the order parameters and the order of the phase transition. Various aspects of second-order phase transitions in crystals occupied him for the next five years [1968SSCom...6..211B, 1969JPSJS..26..167B, 1969PhLA...28..474B, 1969PhLA...28..659B, 1969SSCom...7..331B, 1971PSSBR..43...11B, Wos:A1972m176400003, 1972SSCom..11...39B, 1973AnPhy..76...72B]. The exception is the paper on statistical properties of focal conic textures in liquid crystals [Wos:A1973q279800017], published in 1973 jointly with P.-G. de Gennes and four other authors. It has an interesting mathematical background: the authors describe iterative filling of space with smectic material using the Apollonian packing of circles.
In late 60’s and early 70’s, Nino produced three books dealing with topics in thermodynamics and statistical physics [boccara1968principes, boccara1970physique, boccara1976symetries]. The first of them, Les principes de la thermodynamique classique, is especially worth mentioning because it is a unique book in the statistical physics literature. It presents the “classical” thermodynamics using the axiomatic approach starting from the concept of state. Fundamental thermodynamic quantities are derived as functions of state with some postulated properties, which are then shown to be uniquely defined. The book is as mathematically rigorous as needed and yet it is also very concise, managing to cover all essential topics in only 160 pages. I was pleasantly surprised that it has been recently digitized under the FeniXX program [^2] FeniXX réédition numérique, www.ebooksfenixx.fr , thus it is available in the electronic format from major online bookstores, including Amazon.
In approximately 1976, Nino started studying more abstract solvable models exhibiting phase transitions and related phenomena. These included $n$-vector model [1976PhLA...56..161B], dilute magnetic models [1979PhLA...70..347B, 1981JAP....52.1728B], spin glasses [1981PhLA...86..181B, 1982JAP....53.2192B], as well as various flavours of the Ising model [1982JPhC...15.1381B, Wos:A1983rk79400002]. I believe this happened because of the strong interest in spin glasses and dilute magnetic materials which was developing at that time in the physics community, and also because this was a subject of experimental investigations at Saclay in the solid state division. One of the most interesting and frequently cited papers of this period is the paper from 1977 [Wos:A1977cu51900006], written jointly with two other authors, considering an assembly of asymmetric ellipsoids coupled by a constant infinite-range interactions. This is a model related to the Potts model and it exhibits two phase transitions. The paper demonstrates that the model can be solved exactly.
Almost all papers which followed, from [1982JPhC...15.1381B] to [1989JPCM....1.5721B], written in the period 1982–1989, deal with the Ising model and spin glass models. One of the most important of these is [1983PhLA...94..185B], titled Dilute Ising models: a simple theory. In this paper Nino proposed a conceptually simple and elegant method to explain behaviour of diluted magnetic system models. The method had the simplicity of the mean field approximation but yielding much more accurate results. It was based on the theorem that a function defined on a finite set can be represented by a polynomial.
The papers on Ising model and its various generalizations and modifications form a substantial body of work, with over 20 publications in the period 1982-1989. They are mostly written in collaboration with A. Benyoussef (Faculté des sciences de Rabat, Morocco) and after 1986 also with several other coauthors. Some of these papers investigate very complicated models, for example, in the system considered in [1986PhRvB..34.7775B], the parameter space is six dimensional and the renormalization group transformation which the authors use has 69 fixed points describing a large variety of critical behaviors.
4 Complex systems, 1989-2010
After obtaining the appointment at the University of Illinois at Chicago in 1989, Nino focused his research on cellular automata and models of complex systems. This shift marked a significant change in his research direction, as he did not publish any more papers in statistical physics per se, apart from two papers on the Potts model on fractals [Wos:A1991fe35700004, Wos:A1992gw43300006] where he is listed as a last author. This suggests that Nino wanted to make a significant impact in a different field and was determined to move forward without looking back.
He started with the study of phase transitions in probabilistic cellular automata [PhysRevA.39.3094] and dynamics of “defects” in deterministic rules [bnr90, bnr91, br91]. In 1993, he published an important (in my opinion) paper [b93] which is not as widely known as it should, in which he systematically studied surjective homomorphisms (factors) between elementary rules, focusing on homomorphisms which are themselves CA rules. His hope, I believe, was to find conjugacies between rules, but conjugacy requires bijection and there are no non-trivial elementary rules which would be bijections, so he considered the “next best thing”, namely surjective rules.
Figure 2: Top: UIC Physics Department, 1995, N.B. sitting first from the right. Middle left: N.B. in front of his Mac computer in UIC office, mid 90’s. Middle right: with graduate students at UIC, 1995. Bottom left: author with N.B., graduation ceremony at UIC, 1997. Bottom right: lecture at Automata 2005, Gdańsk. B&W photos from the UIC Physic Dept. brochure (1995), the rest from author’s collection.
Around 1992 he came out with the idea of site exchange cellular automata, that is, cellular automata where site values of selected sites are repeatedly exchanged between iterations. The selection is usually random, but the sites selected for exchange can be either close neighbours (short range exchange) or completely uncorrelated (long range exchanges). He wanted to see how frequent these exchanges need to be to bring the cellular automaton to its mean-field or near mean-field regime. This idea appeared for the first time in the paper written with his first graduate student at UIC, K. Cheong [bc92]. The cellular automaton used in the paper is a model of the epidemic process on a two-dimensional lattice. One of the main findings of this study was that when the rate of site exchanges tends to infinity the correlations are completely destroyed and the time evolution of the epidemic is then correctly predicted by the mean-field approximation.
The results of [bc92] were quickly extended to a more general SIS epidemic model [bc93], in which a transcritical bifurcation similar to a second-order phase transition has been found. The third paper in the series [bco94] added births and deaths to the epidemic model. This time the model revealed oscillatory behaviour of the densities of susceptible and infective individuals as a function of time through a Hopf-type bifurcation parametrized by the rate of site exchanges. A similar phenomenon appears in a version of the predator-prey model studied in [brr94b]. The papers [bc92, bc93, bco94, brr94b] belong to the most frequently cited works of Nino Boccara among his papers on CA and complex systems, and they continue to attract interest even now. Later he co-authored another paper on population models investigating the spread of rabies among foxes [bbcez99].
One of the problems which occupied Nino in early 90’s was the question of existence of collective global behaviour in cellular automata. Following the work of H. Chaté and P. Manneville it was known that in “classic” cellular automata in dimensions lower than four such behaviour is not possible. In [bco94] (published in 1994 but based on results obtained in 1993 or earlier) global oscillations have been observed in two dimensions, but this was because the range of interactions was made effectively infinite by the site exchange process. Nino wondered if such increase of range could be made explicit by considering a weighted totalistic CA in which the weight depends on the distance from the central site $r$ as some function decreasing with $r$, for example, decreasing exponentially or as a power law. He gave this problem to Servie Geurten, a newly arrived graduate student from Holland. I also started to work with Nino as my graduate supervisor almost at the same time (perhaps a few months later), thus we became good friends with Servie. Unfortunately, shortly afterwards a tragedy struck: on December 30, 1993 Servie died in a car accident. We felt devastated but I took over Servie’s notes and we continued working on this problem for a while. Although it eventually drifted into a different direction then originally planned, we published a paper with Servie’s name listed as a coauthor reporting the results. Due to editors sitting on it for a very long long time, it did not appear until 1997 [bfg94].
After the famous paper of K. Nagel and M. Schreckenberg A cellular automaton model for freeway traffic appeared in 1992, interest in CA traffic models was growing rapidly. Nino became interested in this as well, and gave me the paper of Nagel and Schreckenberg to read. We started to look into various aspects of traffic modelling and quickly realized that although some binary CA which conserve the sum of states can be used as simple models of traffic flow, there are many others which are interesting on their own right and these should be studied further. Nino coined the name “number-conserving cellular automata” for them, and we published several papers investigating their properties [paper5, paper8, paper19, paper12, paper28], the last three being the result of continued collaboration after I left UIC with PhD in 1997. He later generalized the idea of number-conserving CA to monotone and eventually-conserving CA [MOREIRA2004285, bocc2007a].
Although I remarked earlier that after 1989 Nino have not published much in the “classical” statistical physics, he still used various methods of statistical physics to study dynamics of cellular automata. For example, the 1993 paper [obs93], written with G. Ódor and G. Szabó, investigates the effect of mixing on one-dimensional probabilistic cellular automaton with totalistic rule, and it uses multiple-point-correlation approximation. The authors found the tricritical point for this system and estimated the critical exponent $\beta$. The paper with F. Bagnoli and P. Palmerini [bbp98] from 1998 studies a different probabilistic rule, namely totalistic cellular automaton having two absorbing states. The automaton is found to exhibit the second-order phase transition with the same critical behavior as the directed percolation model, in qualitative agreement with mean-field predictions.
In the last five years of his active career Nino studied models of the spread of globalized culture and opinion formation [bocc2005, bocc2007, bocc2008, WOS:000263880100006, WOS:000271586900002, WOS:000278088800005, WOS:000275469100001]. He was deeply concerned about what American sociologist George Ritzer called the McDonaldization of society, and wondered what could be done to keep alive diverse local cultures and languages. In [bocc2007] he proposed a model for the spread of a globalized culture in a population of individuals located at the vertices of a scale-free social network. This was a model with synchronous update, thus one could call it a generalized CA model. His conclusions were not very optimistic, but nevertheless recommended some concrete solutions:
In essence, this model shows that to resist globalization and keep alive a diverse local culture, role models are essential, creations of as many links as possible between individuals sharing the same local culture has to be encouraged, and the society as a whole must make every effort to protect and promote its own culture even at the cost of not fully respecting the rules of free trade.
The last paper of Nino Boccara was published in 2010 [WOS:000275469100001], but I believe his last large project was the second edition of the book Modeling Complex Systems [compsys2003]. Originally published in 2003 by Springer, the book served as a textbook with the purpose of instructing readers in the construction of models of complex systems and equipping them with mathematical tools needed to analyze their dynamics. It was quite a success and received glowing reviews in Physics Today, Journal of Statistical Physics and several other leading physics journals and magazines. In 2009, Nino decided to update it and prepared the second edition, adding many new exercises with solution, extensively revising references and footnotes, and adding many features improving its usability as a textbook. The book includes discussion of Nino’s own work, such as site-exchange, number-conserving and generalized cellular automata. It remains in print and I am sure it will serve as a testament of his legacy for many years to come.
Acknowledgements I wish to thank Nazim Fatès and Anna Lawniczak for encouragement to write this article, Franco Bagnoli for contributing his recollections and M. Catherine Kounelis for information about N.B. appointment at ESPCI. I am also very grateful to Nazim Fatès for reading and correcting the manuscript.
5 Bibliography of works of Nino Boccara
{btSect}
[hfbib]bocc-books
5.1 Books
[…]
[hfbib]bocc-edited
5.2 Edited books
[…]
[hfbib]bocc-articles-early-new
5.3 Research papers 1960–1989
[…]
[hfbib]bocc-articles
5.4 Research papers 1989–2010
[…]